# A Strong Baseline for Temporal Video-Text Alignment

### v1

@@ 3.3 Architecture
As shown in Fig. 3, we adopt a simple Transformer-based architecture, where visual features of each clip are individually encoded and treated as key-value pairs, the textual features of narrations or steps are treated as queries. The queries iteratively attend visual features with cross attention, and the alignment scores can be computed by textual-visual similarity, which emits high scores for any alignable sentences with their corresponding video timestamp. The following sections describe the full details.
3.3.1 Visual-Textual Features
@@
As defined in Sec. 3.1, given an untrimmed long instructional video, 8-mins long in our case, with associated texts (narrations or procedural steps), we compute the visual and textual features with pre-trained backbones, that can be denoted as:
@@
where refers to features of a video sequence, and denotes the textual features for narrations or procedural steps associated with the video. is the dimension of the feature vector. Note that, the resulting feature dimensions depend on the pre-trained visual or textual backbone.
@@
We consider three popular pre-trained visual-language models, namely, the S3D-G backbone trained with MIL-NCE [xie2018rethinking, miech2020end], CLIP-ViT/L-14 trained with InfoNCE [oord2018representation, radford2021learning], and InternVideo-MM-L14 trained with masked video reconstruction and multi-modal contrastive loss [wang2022internvideo, wang2023videomae]. We follow the pre-processing procedure from corresponding models, as detailed below:
S3D-G. As for pre-processing, the video is decoded with 16 frames per second (fps). Each 16-frame video clip cropped by a non-overlapping temporal window is fed into the S3D-G architecture, resulting in one feature (512-d) per second. For each sentence, it is encoded with the text encoder associated with the S3D-G video encoder [miech2020end]. Specifically, the sentence is first tokenized and converted into word vectors (300-d) through an embedding layer trained on Google News. Then, word vectors from the same sentence will be projected into 512-d vectors, and turned into one sentence vector through maxpooling.
CLIP. We decode the video into 1fps and extract one visual feature (768-d) per second with OpenAI’s ViT/L-14 model [dosovitskiy2020image]. The text feature (768-d) is encoded with CLIP’s text encoder. The embedding of the <eot> token is used as the feature for each sentence.
InternVideo. We decode the video into 8fps and feed it into the InternVideo-MM-L14 model with an 8-frame non-overlapping temporal window, obtaining one visual feature (768-d) per second. The text encoder is the same as CLIP’s but pre-trained on video-text datasets. Similarly, the embedding of the <eot> token is feature of each sentence.
3.3.2 Video-Text Alignment Module
@@
As shown in Fig. 3, after extracting the visual and textual features independently, we project both features into the same dimension, fuse multimodal information with Transformer, and then predict the alignment score matrix between video and texts.
@@
Feature projection. After computing visual-textual features with pre-trained frozen models, we adopt one linear layer to project the features into the embedding with the same dimension . In terms of positional encoding, we add sin/cos positional encoding to the visual features, while on the text side, we only add learnable positional embeddings to sentence features for the narration alignment task, but not in the step grounding task:
@@
Visual-textual feature fusion. The visual features are processed with a temporal aggregator, followed by a grounding module, expressed as:
@@
Alignment prediction. In order to get the alignment score matrix between video and texts, we first project the output of the encoder and decoder into same dimension,
@@
where is the predicted alignment matrix. The higher value of means the -th sentence is more likely to align with the visual content of timestamp .
Training Details. As introduced in Sec. 3.1, given the ground truth labels , we train the video-text alignment network with a variant of the InfoNCE loss, which has been adopted for such video-text alignment problem [han2022temporal, mavroudi2023learning]:
where is the ground truth alignment matrix. is a temperature hyper-parameter, and refer to -th sentence and -th timestamp in the video respectively.
@@ 4 Experiments
In this section, we start by describing the datasets used in this paper, then present the implementation details and results for the video-text alignment task.
| ASR Transcripts | HT-Step | HTM-Align |
| R@1 | R@1 | |
| Youtube | 36.1 | 60.4 |
| WhisperX | 35.9 | 63.8 |
@@ 4.1 Datasets and Metrics
Following the previous work [han2022temporal, mavroudi2023learning], we train our NSVA model on a subset of the HowTo100M dataset, with narrations from transcripts and steps generated in Sec. 3.2. As for evaluation, we conduct procedural step grounding on HT-Step [mavroudi2023learning], narration alignment on HTM-Align [han2022temporal], and zero-shot action step localization on CrossTask [zhukov2019cross].
HTM-370K (Training). The HowTo100M dataset [miech2019howto100m] is a large-scale instructional dataset crawled from YouTube, consisting of approximately 1.2M videos with ASR transcripts. Following previous work [han2022temporal, mavroudi2023learning], we use HTM-370K for training, it contains videos from the Food & Entertaining categories, consisting of 32% of the videos of the entire HowTo100M dataset.
HowToStep (Training). As introduced in Sec. 3.2, we construct this dataset for training, by transforming the original transcripts of HTM-370K into around M ordered instructional steps with start/end timestamps for almost 340K videos after filtering.
HTM-Align (Evaluation). This benchmark is proposed by [han2022temporal] for evaluating narration alignment. It contains 80 videos from the HTM-370K as a holdout testing set. The authors have manually labeled the alignability for each narration and further aligned them to the visual signal with start/end timestamps if alignable. The metric on this dataset is Recall@1 (R@1), which means if the maximally matched timestamp for each alignable sentence model predicted falls into the ground truth temporal window, it is regarded as being successfully recalled. The recall score is computed as the ratio of successfully recalled sentences to all the alignable sentences.
HT-Step (Evaluation). This benchmark [mavroudi2023learning] aims to evaluate the procedural step grounding. It contains manual annotations for 600 videos, specifically, for each video, the authors first collect activity steps from the related Wikihow article using the task name, e.g., Make Pumpkin Puree, and then annotates the temporal segment for steps alignable with the video. The metric on this dataset is same as HTM-Align, namely, R@1.
CrossTask (Evaluation). In addition to benchmarks based on HowTo100M, we also adopt this established instructional video benchmark for zero-shot step localization. The CrossTask Dataset [zhukov2019cross] contains 4800 videos, which can be divided into 18 primary tasks (cooking, DIY car, and shelf assembly) and 65 related tasks. The videos in the primary tasks are annotated as steps with temporal segments from a predefined taxonomy of 133 steps. The metric on this dataset is Average Recall@1 (Avg. R@1), which measures the recall over steps in videos for each task and averages the results. We take a random set of 1850 videos from 18 primary tasks as the evaluation set as previous work [mavroudi2023learning].
| Backbone | HT-Step | HTM-Align |
| R@1 | R@1 | |
| CLIP ViT/L-14 [radford2021learning] | 33.3 | 53.9 |
| MIL-NCE S3D-G [miech2020end] | 29.1 | 56.4 |
| InternVideo ViT-L/14 [wang2022internvideo] | 35.9 | 63.8 |
@@ 4.2 Implementation Details
Overall, we investigate the effectiveness of three popular pre-trained visual-language models, namely, the S3D-G, CLIP-ViT/L-14, and InternVideo-MM-L14, as described in Sec. 3.3. While exploring other factors, for example, the effect of ASR transcripts, and the effect of incorporating descriptive steps during training, we use InternVideo-MM-L/14 by default, unless specified otherwise.
@@
At training time, the temperature in our loss is 0.07. We use the AdamW [loshchilov2017decoupled] optimizer and train the model with an initial learning rate and cosine decay for 12 epochs. When determining the start/end time for generated steps, the hyper-parameters for the 2-stage are , , which is also discussed in ablation study. When training for narration alignment, we add learnable positional embeddings to the extracted textual features and keep the original order of texts, while training for procedural step grounding, we shuffle texts on the train set, with no positional embeddings being added. Complete implementation details are included in the appendix.
4.3 Results
4.3.1 Ablation Studies
@@
We explore the effects of multiple design choices on narration alignment and procedural step grounding, and all models are evaluated on HTM-Align and HT-Step.
@@
Effect of upgrading ASR system. To start with, we compare the original transcripts generated from YouTube, with that from the recent WhisperX [bain2022whisperx, radford2023robust], as the weakly-aligned labels for training. Qualitatively, we do observe that WhisperX generates fewer punctuation errors, and gives higher accuracy of temporal boundaries in the ASR transcripts. As shown in Tab. 1, upgrading the ASR system indeed leads to noticeable performance improvement in narration alignment. However, in procedural step grounding, it is not effective, showing a marginal performance decrease, we conjecture this is because the gap of text style between the train set (dense spoken narrations) and test set (concise activity steps) dominates the performance. In later sections, we use the WhisperX transcripts by default.
@@
Effect of upgrading visual-textual backbone. Here, we explore different visual backbones with the corresponding text encoder. As shown in Tab. 2, S3D-G exceeds CLIP ViT/L-14 in narration alignment but is inferior to the latter in step grounding. In general, InternVideo ViT/L-14 shows significant advantages on both tasks, attributed to the large pre-trained dataset and effective supervision in video representation learning, which is our default choice.
| Backbone | Text | HT-Step | HTM-Align |
| R@1 | R@1 | ||
| CLIP-ViT/L-14 | W | 33.3 | 53.9 |
| CLIP-ViT/L-14 | W + S | 42.0 | 60.3 |
| S3D-G | W | 29.1 | 56.4 |
| S3D-G | W + S | 43.6 | 62.5 |
| InternVideo-ViT/L-14 | W | 35.9 | 63.8 |
| InternVideo-ViT/L-14 | W + S | 46.7 | 69.9 |
Effect of ASR transformation. We validate the effectiveness of the proposed pipeline to transform noisy ASR transcripts into descriptive steps, with the post-determined temporal segments as an extra training source for video-text alignment. As shown in Tab. 3, on three backbones, we find that the generated dataset is effective for both narration alignment and step grounding. Notably, the average improvement exceeds 10% in HT-Step implies that the generated steps are indeed more descriptive, resembling a similar style as the instructions from the Wikihow knowledge base. For narration alignment, our generated steps add more diversity for training, thus leading to better performance.
Ablation of methods to generate dataset. We investigate the two-stage approach with hyper-parameters to generate the extra training dataset (HowToStep) as described in Sec. 3.2. As shown in Tab. 4, the difference between directly determining the timestamp of the generated steps by video-step similarity matrix and indirectly by chaining ‘stepstranscript’ similarity and ‘transcriptvideo’ labels is not obvious for the first stage, while becomes significant when using the pseudo-label of the first stage for the self-training in the second stage, there is a large gap between the two methods. We find that using the video-step similarity matrix directly will make the model learn a trivial solution (i.e., identity mapping), while indirectly obtained pseudo-label can let the model learn the alignment patterns as analyzed in Sec. 3.2.2. In addition, we choose the best thresholds to generate the final dataset according to Tab. 4.
@@
| Step-Video | 0.15 | ✗ | - | 37.2 |
| Step-Video | 0.15 | ✓ | 0.8 | 37.4 |
| Step-ASR | 0.15 | ✗ | - | 35.1 |
| Step-ASR | 0.35 | ✗ | - | 36.9 |
| Step-ASR | 0.15 | ✓ | 0.7 | 42.4 |
| Step-ASR | 0.15 | ✓ | 0.8 | 43.0 |
| Step-ASR | 0.35 | ✓ | 0.7 | 41.5 |
| Step-ASR | 0.35 | ✓ | 0.8 | 42.5 |
4.3.2 Comparison with State-of-the-art
Based on the conducted ablation studies, we compare our best model with existing state-of-the-art approaches, to present a strong, yet simple baseline on video-text alignment for future research. As shown in Tab. 5, on the challenging HT-Step task, that aims to ground unordered procedural steps in videos, our model achieves 46.7% R@1, leading to an absolute improvement of 9.3%, over the existing state-of-the-art (37.4%) achieved by VINA [mavroudi2023learning]; On HTM-Align [han2022temporal], which aligns narrations in the video, our method exceeds sota model by 3.4%; On CrossTask [zhukov2019cross], where we need to align video frames and task-specific steps without finetuning, our method outperforms existing the state-of-the-art approach by 4.7%, demonstrating our model learns stronger joint video-text representation.
| Method | HT-Step | HTM-Align | CrossTask (ZS) |
| R@1 | R@1 | Avg. R@1 | |
| TAN* [han2022temporal] | 31.2 | 47.1 | - |
| VINA [mavroudi2023learning] | 37.4 | 66.5 | 44.8 |
| Ours | 46.7 | 69.9 | 49.5 |

### v2

@@ 3.2 Multi-Sentence Grounding Model
Overall Architecture. As shown in Fig. 3, we adopt a simple Transformer-based architecture, where visual features of each 1s clip are individually encoded and treated as key-value pairs, the textual features of narrations or steps are treated as queries. The queries iteratively attend visual features with cross attention, and the alignment scores can be computed by textual-visual similarity, which emits high scores for any alignable sentences at their corresponding video timestamp. The following sections describe the full details.
@@
Visual-Textual Features. Given an instructional video with associated texts, we extract the visual and textual features with pre-trained backbones,
@@
where refers to features of a video sequence, and denotes the textual features associated with the video. is the dimension of the feature vector. Note that, the resulting feature dimensions depend on the pre-trained backbone.
@@
We consider three popular pre-trained visual-language models, namely, the S3D-G backbone trained with MIL-NCE [42, 28], CLIP-ViT/L-14 trained with InfoNCE [31, 33, 13], and InternVideo-MM-L14 trained with masked video reconstruction and multi-modal contrastive loss [41, 40]. The ablation studies about these pre-trained backbones are presented in Sec. 4.4.
@@
Multi-Sentence Grounding Module. As shown in Fig. 3, after extracting the visual and textual features independently, we project both features into the same dimension, fuse multimodal information with Transformer, and then predict the alignment score matrix between video and texts.
@@
(1) Feature Projection. After computing visual-textual features with pre-trained frozen models, we adopt one linear layer to project the features into the embedding with the same dimension . In terms of positional encoding, we add sin/cos positional encoding to the visual features:
@@
(2) Visual-Textual Feature Fusion. The visual features are processed with a temporal aggregator, followed by a grounding module, expressed as:
@@
(3) Alignment Prediction. To get the alignment score matrix between video and texts, we project the encoder and decoder outputs into the same dimension,
@@
where is the predicted alignment matrix. The higher value of means the -th sentence is more likely to align with the scene of timestamp .
3.3 Self-training to Refine Generated Steps Grounding
In this section, we describe the procedure to refine the start-end timestamp of generated procedural steps, by first training the proposed multi-sentence grounding model, and use it to update the labels.
Training with Estimated Timestamps of Procedural Steps. Given the approximately estimated timestamps for the generated steps in Sec. 3.1, denoted as , we train the multi-sentence grounding network with a variant of the InfoNCE loss, following [17, 27]:
where is the model’s output alignment matrix, as explained in Sec. 3.2, is a temperature hyper-parameter, and refer to -th sentence and -th timestamp in the video respectively.
Updating the Timestamps of Procedural Steps. We use the trained model to do inference on the whole set of procedural steps generated by LLM. Specifically, when feeding video and a set of descriptive steps as model input, the proposed NaSVA model outputs the alignment matrix (). We take the timestamp of the maximal alignment score as the start time for each step, and the duration is a constant , following [9, 37]. Similarly, the generated step with the maximal alignment score lower than a threshold will be discarded. Finally, we obtain a dataset consisting of aligned descriptive steps, named HowToStep. In practice, we observe that although our NaSVA model is trained on noisy pseudo-labels (i.e., relying on weakly-aligned transcripts, and getting steps-transcript similarity from imperfect pre-trained text encoder), the model tends to learn the alignment patterns from procedural step to videos before overfitting to the noises, as also being observed in [44]. Note this refined process (i.e., self-training) can be repeated multiple rounds. To balance computation cost and performance improvement, we choose only one round as default. More details for generating the dataset are presented in ablation studies in Sec. 4.4 and Appendix as well.
Relation to Existing Work. We note that recent works [10, 27, 2] aim to ground steps collected from the only task-related articles in the external knowledge base (i.e., Wikihow) within videos from HowTo100M as shown in Fig. 1 (middle), while we try to ground the procedural steps summarized from the ASR transcript of the same video. This motivation has two natural advantages: (i) the procedural steps generated per task are more diverse than those from Wikihow (e.g., 74 different videos are showing how to make pumpkin puree in HowTo100M, while only one article of the same task is collected in Wikihow); (ii) the generated steps within the video are more likely to be alignable to the video itself than those externally sourced Wikihow steps intrinsically.
3.4 Downstream Applications
To demonstrate the effectiveness of the automatic pipeline for curating high-quality video-text datasets at scale, we consider using HowToStep for training and evaluation on multi-sentence grounding tasks. As shown in Fig. 4, we consider two variants, depending on the type of given texts (narrations or procedural steps), and differentiate the two targeted tasks as follows:
Narration Alignment [17]. arrations () are textual sentences transcribed from the instructor’s speech via the ASR system (around 100 sentences per video in HTM-Align [17]), which means these sentences maintain a time order that can be utilized for visual grounding. Note that, some narrations are not visually alignable or not well-aligned with visual signals temporally as shown in Fig. 1 (left). However, the ASR transcripts with timestamps can be used as weakly-aligned labels for training multi-sentence grounding network as previous work does [17].
Procedural Step Grounding [27, 48]. Procedural teps () are collected from Wikihow [18], an external knowledge base with instructional articles explaining the procedures for completing certain tasks (around 10 steps per task in HT-Step [27]) or pre-defined action taxonomy (around 5 steps per video in CrossTask [48]). Compared with narrations, the procedural steps do not necessarily have consistent temporal order as that happens in the video, and some steps may not even be present in the video at all.
@@ 4 Experiments
Here, we start by describing the datasets for training and evaluation, then present the implementation details and results for the multi-sentence grounding tasks.
@@ 4.1 Datasets and Metrics
Following the previous work [17, 27], we perform dataset curation on a subset of the HowTo100M dataset, and train our NaSVA model with narrations from transcripts and steps from HowToStep. As for evaluation, we conduct procedural step grounding on HT-Step [27], narration alignment on HTM-Align [17], and zero-shot action step localization on CrossTask [48].
HTM-370K (Training). The HowTo100M dataset [29] is a large-scale instructional dataset crawled from YouTube, consisting of approximately 1.2M videos with ASR transcripts. Following previous work [17, 27], we use HTM-370K for training, it contains videos from the Food & Entertaining categories, consisting of 32% of the videos of the entire HowTo100M dataset.
HowToStep (Training). As introduced in Sec. 3, we construct this dataset for training, by transforming the original transcripts of HTM-370K into around M ordered instructional steps with post-determined start/end timestamps for almost 340K videos after filtering.
HTM-Align (Evaluation). This benchmark is proposed by [17] for evaluating narration alignment. It contains 80 videos from the HTM-370K as a holdout testing set. The authors have manually labelled the alignability for each narration and further aligned them to the visual signal with start/end timestamps if alignable. The metric on this dataset is Recall@1 (R@1), which means if the maximally matched timestamp for each alignable sentence model predicted falls into the ground truth temporal window, it is regarded as being successfully recalled. The recall score is computed as the ratio of successfully recalled sentences to all the alignable sentences.
HT-Step (Evaluation). This benchmark [27] aims to evaluate the procedural step grounding. It contains manual annotations for 600 videos, specifically, for each video, the authors first collect activity steps from the related Wikihow article using the task name, e.g., Make Pumpkin Puree, and then annotate the temporal segment for steps alignable with the video. The metric on this dataset is the same as HTM-Align, namely, R@1.
CrossTask (Evaluation). In addition to benchmarks based on HowTo100M, we also adopt this established instructional video benchmark for zero-shot step localization. The CrossTask Dataset [48] contains 4800 videos, which can be divided into 18 primary tasks and 65 related tasks. The videos in the primary tasks are annotated as steps with temporal segments from a predefined taxonomy. The metric is Average Recall@1 (Avg. R@1), which measures the recall over steps in videos for each task and averages the results. Following previous work [27], we evaluate on a random set of 1850 videos from the primary tasks.
Method | HT-Step | HTM-Align | CrossTask (ZS) |
R@1 | R@1 | Avg. R@1 | |
Zhukov [48] | - | - | 40.5 |
UniVL [26] | - | - | 42.0 |
TAN* [17] | 31.2 | 47.1 | - |
VINA [27] | 37.4 | 66.5 | 44.8 |
Ours | 46.4 | 71.6 | 46.7 |
@@ 4.2 Implementation Details
Overall, we investigate the effectiveness of three popular pre-trained visual-language models for constructing the whole pipeline, namely, the S3D-G, CLIP-ViT/L-14, and InternVideo-MM-L14, as described in Sec. 3.2. While exploring other factors, for example, the effect of ASR transcripts, and the effect of incorporating descriptive steps during training, we use InternVideo-MM-L/14 by default, unless specified otherwise. At training time, the temperature in our loss is 0.07. We use the AdamW [25] optimizer and train the model with an initial learning rate and cosine decay for 12 epochs. When determining the start/end time for generated steps, the hyper-parameters for the 2-stage are , , which are also discussed in the ablation study. We train one unified model for both narration alignment and procedural step grounding tasks, by setting the texts of one training batch to be either ordered, dense narrations or shuffled, sparse procedural steps. Complete implementation details are included in the Appendix.
4.3 Main Results
Comparison with State-of-the-art. We compare our best model with existing state-of-the-art approaches on three public benchmarks for multi-sentence grounding tasks. As shown in Tab. 1, on the challenging HT-Step task, that aims to ground unordered procedural steps in videos, our model achieves 46.4% R@1, leading to an absolute improvement of 9.0%, over the existing state-of-the-art (37.4%) achieved by VINA [27]. On HTM-Align [17], which aligns narrations in the video, our method exceeds the SOTA model by 5.1%. On CrossTask [48], where we need to align video frames and task-specific steps without finetuning, our method outperforms existing the state-of-the-art approach by 1.9%, demonstrating the effectiveness of the proposed pipeline for downstream tasks.
4.4 Ablation Study
@@
We explore the effects of multiple design choices in the proposed pipeline and evaluate them on both narration alignment and procedural step grounding tasks.
@@
Effect of Upgrading ASR System. To start with, we compare the original transcripts scrawled from YouTube, with that from the recent WhisperX [3, 34], as the weakly-aligned labels for training. Qualitatively, we do observe that WhisperX generates fewer punctuation errors, and gives higher accuracy of temporal boundaries in the ASR transcripts. As shown in Tab. 4, upgrading the ASR system indeed leads to noticeable performance improvement in narration alignment. However, in procedural step grounding, showing a marginal performance decrease, we conjecture this is because the gap of text style between the train set (dense speeches) and test set (sparse procedural steps) dominates the performance. In later sections, we use the WhisperX transcripts by default.
@@
Effect of Upgrading Visual-Textual Backbone. Here, we explore different visual backbones with the corresponding text encoder. As shown in Tab. 4, S3D-G exceeds CLIP ViT/L-14 in narration alignment but is inferior to the latter in step grounding. In general, InternVideo ViT/L-14 shows significant advantages on both tasks, attributed to the large pre-trained dataset and effective supervision in video representation learning, which is our default choice in the pipeline.
@@
Step-Video | 0.20 | ✗ | - | 36.5 |
Step-Video | 0.20 | ✓ | 0.8 | 36.9 |
Step-ASR | 0.15 | ✗ | - | 35.3 |
Step-ASR | 0.20 | ✗ | - | 36.0 |
Step-ASR | 0.15 | ✓ | 0.7 | 41.4 |
Step-ASR | 0.15 | ✓ | 0.8 | 41.7 |
Step-ASR | 0.20 | ✓ | 0.7 | 42.3 |
Step-ASR | 0.20 | ✓ | 0.8 | 43.7 |
Effect of the Proposed Dataset. We validate the effectiveness of using this pipeline to transform noisy ASR transcripts into descriptive steps, with the post-determined temporal segments as an extra training source for multi-sentence grounding. As shown in Tab. 2, on all three backbones, the generated dataset is effective for both narration alignment and step grounding. Notably, the average improvement exceeds 10% in HT-Step implies that the generated dataset are indeed more descriptive and well-aligned. For narration alignment, our generated steps add more diversity for training, thus leading to better performance.
Ablation of Options in Constructing the Pipeline. We investigate the choices and hyper-parameters to generate the extra training dataset (HowToStep) described in Sec. 3. As shown in Tab. 3, the difference in determining the timestamp of the generated steps by directly computing the video-step similarity matrix, and indirectly by chaining ‘stepstranscript’ similarity and ‘transcriptvideo’ labels is not obvious for the first stage, while becomes significant when using the pseudo-label of the first stage for the self-training in the second stage, there is a large gap between the two methods. We find that using the video-step similarity matrix directly will make the model learn a trivial solution (i.e., identity mapping), while indirectly obtained pseudo-label can let the model learn the alignment patterns as analyzed in Sec. 3.3. In addition, we choose the best thresholds to generate the final dataset according to Tab. 3.
Backbone | ASR System | HT-Step | HTM-Align |
R@1 | R@1 | ||
CLIP | WhisperX | 32.2 | 58.7 |
S3D | WhisperX | 29.1 | 59.5 |
InternVideo | WhisperX | 34.7 | 70.0 |
InternVideo | Youtube | 35.7 | 61.3 |
Dataset | Alignable | Well-aligned |
HTM-370K | 30.1 | 21.9 |
HowToStep | 60.6 | 52.5 |
4.5 Manual Check
To evaluate the quality of our proposed dataset, we have randomly sampled 10 videos with a total of 853 sentences for a manual check, focusing on the proportion of steps in the dataset that are (i) visually alignable, (ii) and well-aligned with correct temporal boundaries. We compare the results before and after using our pipeline, as shown in Tab. 5, demonstrating the significant quality improvement.